OpenStudy (anonymous):

What is the factored form of 3xsquared -9x+12

jhonyy9 (jhonyy9):

x_1,_2=(9+/- sqrt(81-144))/6=(9+/- sqrt(-63))/6=(9+/- i3sqrt(7))/6= =3(3+/- isqrt(7))/6=(3+/- isqrt(7))/2 - those are solutions

OpenStudy (anonymous):

3x^2 - 9x + 12 =3(x^2-3x+4) =3x(x-3) + 12 it can't be completely factored though. Please make sure the question is posted correctly.

OpenStudy (anonymous):

unless it's acceptable to have complex numbers in the solution.